# Optimizing Parameter Estimation for Electrochemical Battery Model: A Comparative Analysis of Operating Profiles on Computational Efficiency and Accuracy

### v2

@@ Abstract
[inst3]organization=Institute of Physical Chemistry, RWTH Aachen University,city= Aachen, postcode=2074, country=Germany
@@ 1 Introduction
More recently, machine learning models have emerged as a promising alternative, offering a data-driven approach to battery modeling [12]. These models, including neural networks, support vector machines, Gaussian process regression, among others, can capture complex nonlinear relationships in battery behavior without requiring detailed knowledge of underlying physicochemical processes [7]. Machine learning models have shown particular success in predicting battery state of charge, remaining useful life, and capacity degradation patterns [13, 14]. However, their accuracy heavily depends on the quality and quantity of training data, and they may struggle to extrapolate beyond their training range.
@@
Electrochemical models, particularly the Pseudo-Two-Dimensional (P2D) model [15] and its variants, offer a more rigorous approach based on fundamental electrochemical principles. These models excel in providing detailed insights into internal battery states and demonstrate superior prediction accuracy [2]. The P2D model, however, requires solving complex systems of partial differential equations, resulting in significant computational overhead. To address this limitation, researchers developed the Single Particle Model (SPM), which simplifies the P2D framework by representing each electrode as a single representative particle and neglecting electrolyte dynamics [16]. While this simplification substantially reduces computational complexity, it compromises accuracy, particularly under high-current operations where electrolyte effects become significant. This led to the development of the Single Particle Model with Electrolyte Dynamics (SPMe), which maintains some computational advantages while incorporating crucial electrolyte effects [17].
@@
However, the implementation of electrochemical models presents significant challenges, particularly in terms of computational complexity and parameter estimation. The parameter identification process is especially demanding because many crucial parameters cannot be directly measured through experimental methods, necessitating the use of sophisticated optimization algorithms for their estimation. The high-dimensional parameter space of electrochemical models, coupled with the interdependence of these parameters, results in computationally intensive estimation processes that can require substantial time and computational resources. Moreover, electrochemical models typically exhibit overparameterization, a phenomenon where distinct parameter sets can produce identical model outputs, leading to challenges in parameter identifiability and uniqueness of solutions [18].
@@
In addressing these challenges, several parameter estimation methodologies have been widely adopted in the field. Particle Swarm Optimization (PSO) [19, 20], Least Squares methods [21], and Genetic Algorithms (GA) [22, 23] represent the predominant approaches for parameter identification in electrochemical battery models. Our previous research conducted a comprehensive comparative analysis of these three methodologies in the context of electrochemical model parameter identification [24]. The findings demonstrated that while PSO exhibited high accuracy and robustness compared to other methods, it required significantly longer computation times.
@@
The selection of operating profiles for parameter estimation plays a crucial role in the accuracy and efficiency of battery parameter identification. Existing research has primarily relied on standardized battery characterization tests, including Constant-Current-Constant-Voltage (CCCV) charging/discharging at various C-rates, pulse tests, and dynamic stress tests (DST). A comprehensive review of the literature reveals diverse approaches to profile selection for parameter estimation [25]. Several researchers have explored the testing protocols. Jokar et al. employed Constant-Current discharge data at multiple C-rates (C/10, C/2, 1C, 2C, 5C) for parameter estimation [26]. Speltino et al. utilized a combination of constant-current (CC) discharge, charge, and pulse tests [27]. Zhang et al. integrated CCCV tests with DST profiles [28]. Fan developed a novel hybrid approach combining CC discharge (SOC 100-50%) with dynamic profiles (SOC 50-0%), though this non-standardized protocol requires specific experimental setup [20]. Zhang et al. combined Electrochemical Impedance Spectroscopy (EIS) data with CCCV profiles, although EIS testing requires additional specialized equipment [29]. Namor et al. proposed a parameter grouping strategy based on physical significance, utilizing different test profiles for each group. For instance, low-rate CC discharge data were used to estimate capacity-related parameters [30]. Some researchers also have incorporated parameter sensitivity analysis into the identification process. Park et al. categorized battery parameters into five groups based on sensitivity analysis, subsequently employing different operating profiles for each parameter group [31, 32]. However, these grouped estimation approaches often require multiple iterations of parameter estimation algorithms, resulting in extended computation times.
Theoretically, utilizing more diverse testing profiles should yield more accurate parameter estimates. However, this approach significantly increases the computational burden of the estimation process. Moreover, incorporating more testing profiles requires additional time for battery testing, which further extends the overall duration of the process. Currently, there is a notable research gap in identifying which combinations of common battery testing profiles can maintain estimation accuracy while minimizing computational time.
To address this gap, our study examines five standard battery testing protocols: CCCV charging/discharging at various rates (C/5, C/2, 1C) and dynamic profiles (pulse charging/discharging, DST). By generating 31 different combinations of these five basic profiles, we conduct a comprehensive comparative analysis of estimation accuracy and computational efficiency across these combinations. This systematic approach aims to identify the optimal profile combination that ensures accuracy while minimizing estimation time, thereby enhancing the efficiency of battery parameter estimation processes.
@@ 2.1 Generation of Basic Profiles
This section delineates the SPM employed in this investigation and elaborates on the basic operational conditions generated for the study.
@@ 2.1.1 SPM
This study is based on an SPM (see Figure 2). The model describes the lithium-ion concentration dynamics in battery electrodes through a system of partial differential equations.
@@
Equations (1)–(8) constitute the complete set of governing equations for the SPM model. Upon examination, we observe that the term appears consistently throughout these equations. This observation leads to an important insight regarding parameter estimation: rather than separately estimating and , we can estimate their combined form as a single parameter. This approach, known as parameter grouping, effectively reduces the number of parameters requiring estimation while maintaining model accuracy. Moreover, each parameter cannot be uniquely identified independently.
@@
where and are the initial concentrations of lithium in the solid phase for the positive and negative electrodes, respectively. Through the implementation of parameter grouping methodology, the parameter estimation process is significantly streamlined, requiring only the nine parameters defined in Equation (12). This reduced parameter set maintains the model’s fidelity while substantially decreasing the computational complexity of the estimation procedure.
@@ 2.1.2 Basic Profiles
In this study, a commercial 18650 lithium-ion battery was employed as the experimental subject. The battery features a nickel manganese cobalt oxide (NMC) cathode and a graphite anode, with a nominal capacity of 2.9 Ah. Parameter identification has been previously conducted on this battery, yielding a set of battery model parameters as presented in Table 1. These identified parameters are utilized as the true parameters of the battery model. Based on these parameters and the established battery model, we generated data for five typical operating conditions, obtaining current and voltage values to facilitate subsequent battery parameter identification procedures.
@@
Based on the SPM model described in the previous section, five basic operating conditions were generated for battery parameter estimation: CCCV discharge-charge at C/5, C/2, and 1C rates, pulse discharge and charge profiles, and DST. Figure 3 illustrates the current inputs for these five basic operating conditions. The experimental protocols detailed in Table 2 present a comprehensive matrix of battery testing conditions.
@@ 2.2 Test Matrix Development
In this section, we generate multiple composite operating conditions by combining the five basic conditions obtained in the previous section. The CCCV and pulse profiles inherently contain both discharge and charge processes, with each cycle beginning and ending at a fully charged state, allowing for direct sequential connection of these profiles. A notable consideration pertains to the DST profile, which exclusively comprises discharge operations without corresponding charge processes. Moreover, the DST protocol is specifically designed to commence from a fully charged battery state. Consequently, when constructing composite operating conditions that incorporate the DST profile, it must be positioned as the terminal segment in the sequence to maintain operational consistency and ensure proper initial conditions are met for each test component.
@@ 2.3 Parameter Estimation and Evaluation
The 31 operating conditions derived in the previous section serve as input profiles for parameter estimation. This study employs the PSO algorithm for parameter identification. The PSO method has been widely applied to battery parameter estimation [2]. Furthermore, building upon our previous research which demonstrated PSO’s effectiveness as a global search algorithm with notable accuracy and stability characteristics [24]. All computational implementations in this study were developed in Python programming language. The PSO algorithm was implemented using the ’pyswarms’ library (version 1.3.0). Parameter estimation was performed on a computing system equipped with a 16-core CPU processor, 32GB of RAM, running a Linux operating system. The optimization process can be expressed mathematically as:
@@
To evaluate the accuracy of each parameter set, we employ a comprehensive validation approach using the Root Mean Square Error (RMSE) metric. Our validation approach systematically assessed each of the 31 parameter sets by simulating all 31 operating conditions and calculating their respective RMSE values. This generated a total of 961 validation test (31 parameter sets × 31 operating conditions), providing a thorough assessment of how each parameter set performed across the complete range of operational conditions. This extensive cross-validation methodology enabled us to rigorously evaluate the robustness and generalization capability of each parameter set across diverse operational Scenarios. The RMSE for each parameter set is calculated as:
@@ 3 Results & Discussion
The obtained results are now presented in terms of different simulations and metrics. First, we show the case where simulations were performed for each of the five considered charge/discharge profiles (i.e. C5, C2, 1C, P and DST) as well as all possible combinations of these profiles, giving rise to the 31 profiles tested here. First, each of these 31 profiles are used in an optimization problem to find the optimal parameters for that particular profile by using a global minimizer as explained above. Thus, 31 optimal parameter sets in (14) are obtained, one per used profile, which is denoted as Cases. Each of these parameter vectors is then evaluated in terms of output voltage error, the RMSE in (15), for all the 31 considered profiles, which is denoted as Scenarios. A total of 961 validation test (31 parameter sets × 31 operating conditions) can be obtained. This yields the 3131 matrix shown in Figure 4 whose x-axis corresponds to the Scenarios (31 operating conditions) and the y-axis corresponds to the Cases (31 parameter sets). Each row of this matrix reflects a different parameter estimate for each considered profile, and each column is the model evaluation of that particular parameter vector considering the specific profile.
@@ 3.1 L0: Validation by Scenario 1
Analysis of both Figure 5 and Table 4 reveals that Case 1 required the highest total time among all Cases. The total time is calculated by Equation (19). This extended duration is a direct consequence of its comprehensive inclusion of all five operating conditions, resulting in prolonged experimental procedures and parameter estimation processes. A particularly noteworthy observation is that Case 21{C/2,1C}’s parameter estimation error was second only to Case 1, while requiring merely 14% of Case 1’s total time. In other words, Case 21 achieved parameter estimation approximately seven times faster than Case 1.
@@ 3.2 L1: Overall Minimum RMSE Value
The L1 level is found by inspecting the minimum value(0.0073V) of the RMSE corresponding to the position (1,31) in the matrix using the (Case,Scenario) notation. In other words, use the first parameter set obtained from{C/5, C/2, 1C, P, DST} and verify it with the 31st operating condition {DST}. This tells us that the overall minimum error is obtained when the complete dataset is used to estimate the model parameters and only the DST data is used for model validation. Thus, the error obtained by considering the full dataset is larger than the one resulting from using DST alone (position (1,1) in the matrix), which implies that the error associated to the complementary part of the DST data in the full dataset (where full dataset is {C/5, C/2, 1C, P, DST} and the complement of the DST in the full dataset is {C5, C2, 1C, P}) involves a larger error than the DST itself. This statement is verified through the following derivation

### v3

@@ Abstract
[inst3]organization=Institute for Materials Research (IMO-imomec), Hasselt University,addressline=Martelarenlaan 42,city=Hasselt,postcode=3500,country=Belgium
[inst4]organization=Institute of Physical Chemistry, RWTH Aachen University,city= Aachen, postcode=2074, country=Germany
@@ 1 Introduction
More recently, machine learning models have emerged as a promising alternative, offering a data-driven approach to battery modeling [12]. These models, including neural networks, support vector machines, Gaussian process regression, among others, can capture complex nonlinear relationships in battery behavior without requiring detailed knowledge of underlying physicochemical processes [7]. Machine learning models have shown particular success in predicting battery State of Charge(SOC), remaining useful life, and capacity degradation patterns [13, 14]. However, their accuracy heavily depends on the quality and quantity of training data, and they may struggle to extrapolate beyond their training range.
@@
Electrochemical models, particularly the Pseudo-Two-Dimensional (P2D) model [15] and its variants, offer a more rigorous approach based on fundamental electrochemical principles. These models excel in providing detailed insights into internal battery states and demonstrate superior prediction accuracy [2]. The P2D model, however, requires solving complex systems of partial differential equations, resulting in significant computational overhead. To address this limitation, researchers developed the Single Particle Model (SPM), which simplifies the P2D framework by representing each electrode as a single representative particle and neglecting electrolyte dynamics [16]. While this simplification substantially reduces computational complexity, it compromises accuracy, particularly under high-current operations where electrolyte effects become significant. This led to the development of the Single Particle Model with Electrolyte Dynamics (SPMe), which maintains some computational advantages while incorporating crucial electrolyte effects [17]. Nevertheless, the SPM has been widely adopted owing to its computational efficiency and structural simplicity.
@@
However, the implementation of electrochemical models presents significant challenges, particularly in terms of computational complexity and parameter estimation. The parameter identification process is especially demanding because many crucial parameters cannot be directly measured through experimental methods, necessitating the use of sophisticated optimization algorithms for their estimation. The high-dimensional parameter space of electrochemical models, coupled with the interdependence of these parameters, results in computationally intensive estimation processes that can require substantial time and computational resources. Moreover, electrochemical models typically exhibit overparameterization, a phenomenon where distinct parameter sets can produce identical model outputs, leading to challenges in parameter identifiability and uniqueness of solutions [18, 19].
@@
In addressing these challenges, several parameter estimation methodologies have been widely adopted in the field. Particle Swarm Optimization (PSO) [20, 21], Least Squares methods [22], and Genetic Algorithms (GA) [23, 24] represent the predominant approaches for parameter identification in electrochemical battery models. Our previous research conducted a comprehensive comparative analysis of these three methodologies in the context of electrochemical model parameter identification [25]. The findings demonstrated that while PSO exhibited high accuracy and robustness compared to other methods. Beyond the three mainstream approaches, a variety of alternative metaheuristic algorithms have also been applied to electrochemical model parameter estimation. For instance, Tian et al. introduced the Two-Population Grey Wolf Optimization (TPGWO) for P2D model identification [26]. Huang et al. enhanced the standard PSO with a Cuckoo Search strategy, proposing a multi-step Cuckoo Particle Swarm Optimization (MCPSO) to accelerate convergence in SPMe parameterization [27]. Tim at al. proposed a strategically switched metaheuristic (SSM) framework, where a convolutional neural network dynamically recommends the most suitable metaheuristic and resampling is used to avoid local optima [28]. Nevertheless, PSO is commonly adopted as a benchmark method for comparison in such studies.
Parameter identification strategies can further be categorized into single-shot identification and stepwise identification [29]. In single-shot identification, all parameters are optimized simultaneously using the full dataset and a global objective function, ensuring consistency and reducing the complexity of variable dependencies. In contrast, stepwise identification decomposes the problem into multiple sequential stages, often starting with parameters that are more easily observable (e.g., ohmic resistance or capacity) and subsequently refining the more complex kinetic and diffusion-related parameters [30]. While stepwise identification can sometimes improve convergence and interpretability, it typically requires multiple runs of the estimation algorithm, parameter sensitivity analysis, and careful control of operating conditions at each stage, which substantially increases computational cost and experimental complexity.
@@
The selection of operating profiles for parameter estimation plays a crucial role in the accuracy and efficiency of battery parameter identification. Existing research has primarily relied on standardized battery characterization tests, including Constant-Current-Constant-Voltage (CCCV) charging/discharging at various C-rates, pulse tests, and dynamic stress tests (DST). A comprehensive review of the literature reveals diverse approaches to profile selection for parameter estimation [29]. Several researchers have explored the testing protocols. Jokar et al. employed Constant-Current discharge data at multiple C-rates (C/10, C/2, 1C, 2C, 5C) for parameter estimation [31]. Speltino et al. utilized a combination of constant-current (CC) discharge, charge, and pulse tests [32]. Zhang et al. integrated CCCV tests with DST profiles [33]. Fan developed a novel hybrid approach combining CC discharge (SOC 100-50%) with dynamic profiles (SOC 50-0%), though this non-standardized protocol requires specific experimental setup [21]. Zhang et al. combined Electrochemical Impedance Spectroscopy (EIS) data with CCCV profiles, although EIS testing requires additional specialized equipment [34]. Namor et al. proposed a parameter grouping strategy based on physical significance, utilizing different test profiles for each group. For instance, low-rate CC discharge data were used to estimate capacity-related parameters [35]. Some researchers also have incorporated parameter sensitivity analysis into the identification process. Park et al. categorized battery parameters into five groups based on sensitivity analysis, subsequently employing different operating profiles for each parameter group [36, 37]. However, these grouped estimation approaches often require multiple iterations of parameter estimation algorithms, resulting in extended computation times. Chun and Han employed a reinforcement learning approach to design specific current profiles for battery parameter estimation [38]; however, this method requires separate experimental tests under specially designed operating conditions.
Theoretically, employing a greater variety of testing profiles can improve the accuracy of parameter estimation. However, such an approach inevitably increases both the computational burden and the experimental effort required for data acquisition, thereby extending the overall estimation process. Existing studies have not systematically addressed the trade-off between accuracy and efficiency, leaving a research gap in determining which combinations of commonly used testing profiles can achieve reliable accuracy while minimizing computational cost.
To bridge this gap, the present study provides one of the first systematic evaluations of testing profile combinations for electrochemical model parameter estimation. Specifically, we investigate five widely adopted battery testing protocols—CCCV charging/discharging at different rates (C/5, C/2, 1C) and dynamic profiles (pulse charging/discharging, DST). By generating 31 distinct combinations of these protocols, we conduct a comprehensive comparative analysis to quantify both estimation accuracy and computational efficiency. It is worth emphasizing that all testing profiles employed in this study are standard laboratory procedures routinely used for model parameterization and validation in both research and industrial battery testing. These protocols effectively capture the essential behaviors observed under practical charging, discharging, and dynamic operating conditions, ensuring that the findings are highly relevant and directly applicable to real-world parameter estimation workflows. This systematic approach enables the identification of optimal profile combinations that ensure robust accuracy with reduced estimation time, thereby providing practical guidance for accelerating electrochemical model development and deployment in battery management applications.
The structure of this paper is as follows. Section 2 introduces the methodology, including the SPM and the experimental design covering 31 testing profile combinations. Section 3 presents and discusses the results, highlighting the trade-offs between estimation accuracy and computational efficiency, and identifying optimal profile combinations for different applications. Section 4 concludes the study by summarizing the key contributions and practical implications for battery state estimation.
@@ 2.1 Generation of Basic Profiles
Given its computational efficiency and simpler structure relative to the P2D model, the SPM has been extensively studied and applied in engineering practice [2]; thus, it is selected in this study.
@@ 2.1.1 SPM
This study is based on an SPM (see Figure 2). The model used in this study is derived from our previous research, and further details can be found in Reference [39, 40]. The model describes the lithium-ion concentration dynamics in battery electrodes through a system of partial differential equations.
@@
Eq. (1)–(8) constitute the complete set of governing equations for the SPM model. Upon examination, we observe that the term appears consistently throughout these equations. This observation leads to an important insight regarding parameter estimation: rather than separately estimating and , we can estimate their combined form as a single parameter. This approach, known as parameter grouping, effectively reduces the number of parameters requiring estimation while maintaining model accuracy. Moreover, each parameter cannot be uniquely identified independently.
@@
where and are the initial concentrations of lithium in the solid phase for the positive and negative electrodes, respectively. Through the implementation of parameter grouping methodology, the parameter estimation process is significantly streamlined, requiring only the nine parameters defined in Eq. (12). This reduced parameter set maintains the model’s fidelity while substantially decreasing the computational complexity of the estimation procedure.
@@ 2.1.2 Basic Profiles
In this study, a commercial 18650 lithium-ion battery was employed as the experimental subject. The battery features a nickel manganese cobalt oxide (NMC) cathode and a graphite anode, with a nominal capacity of 2.9 Ah. Parameter identification has been previously conducted on this battery, yielding a set of battery model parameters as presented in Table 1.The data in Table 1 are derived from our previous study, where parameter estimation was performed using PSO on real battery testing data. For further details, please refer to Reference [39]. These identified parameters are utilized as the true parameters of the battery model. Based on these parameters and the established battery model, we generated data for five typical operating conditions, obtaining current and voltage values to facilitate subsequent battery parameter identification procedures.
@@
Based on the SPM model described in the previous section, five basic operating conditions were generated for battery parameter estimation: CCCV discharge-charge at C/5, C/2, and 1C rates, pulse discharge and charge profiles, and DST. Figure 3 illustrates the current inputs for these five basic operating conditions. The experimental protocols detailed in Table 2 present a comprehensive matrix of battery testing conditions. Since the SPM neglects concentration variations in the electrolyte, its prediction errors typically increase under high current conditions, generally above 1C [42]. As shown in the second column of Table 2, the maximum equivalent discharge rate across all five operating profiles is C/1.4, which is below 1C. Although the DST profile includes brief 2C discharge pulses, it does not involve long-duration 2C discharges. Therefore, the operating conditions considered in this study remain consistent with the applicability of the SPM.
@@ 2.2 Test Matrix Development
In this section, we generate multiple composite operating conditions by combining the five basic conditions obtained in the previous section. The CCCV and pulse profiles inherently contain both discharge and charge processes, with each cycle beginning and ending at a fully charged state (SOC = 1). This characteristic allows these profiles to be directly connected in sequence without introducing inconsistencies in the SOC. A notable consideration pertains to the DST profile, which exclusively comprises discharge operations starting from SOC = 1 and terminating at SOC = 0. Furthermore, the DST protocol is specifically designed to initiate from a fully charged battery state. For this reason, the DST must be positioned as the terminal segment in any composite sequence to ensure that the required initial conditions are satisfied and that no subsequent profiles are compromised by its terminating at a depleted state.
It is also important to clarify that the sequence of the operating conditions does not affect the parameter estimation results in our study. While the DST must remain last, all other profiles begin and end at SOC = 1, which guarantees that the starting point of the following profile remains unaffected regardless of the order in which they are arranged. Consequently, these profiles can be reordered freely without influencing the subsequent states of charge. Moreover, in the PSO-based optimization framework adopted in this work, the loss function is defined as the root-mean-square error (RMSE) over the entire voltage sequence. Because all non-DST profiles consistently return to SOC = 1, reordering them does not alter the overall RMSE calculation and therefore does not influence the optimization process.
@@ 2.3 Parameter Estimation and Evaluation
The 31 operating conditions derived in the previous section serve as input profiles for parameter estimation. In this study, the PSO algorithm is employed for parameter identification. PSO has been widely applied to battery parameter estimation and has demonstrated strong global search capability, accuracy, and stability in our previous work [2, 25]. Therefore, it is adopted here as the optimization method for parameter estimation of the battery model. All computational implementations in this study were developed in Python programming language. The PSO algorithm was implemented using the ’pyswarms’ library (version 1.3.0). Parameter estimation was performed on a computing system equipped with a 16-core CPU processor, 32GB of RAM, running a Linux operating system. The optimization process can be expressed mathematically as:
@@
To evaluate the accuracy of each parameter set, we employ a comprehensive validation approach using the RMSE metric. Our validation approach systematically assessed each of the 31 parameter sets by simulating all 31 operating conditions and calculating their respective RMSE values. This generated a total of 961 validation test (31 parameter sets × 31 operating conditions), providing a thorough assessment of how each parameter set performed across the complete range of operational conditions. This extensive cross-validation methodology enabled us to rigorously evaluate the robustness and generalization capability of each parameter set across diverse operational Scenarios. The RMSE for each parameter set is calculated as:
@@ 3 Results & Discussion
The obtained results are now presented in terms of different simulations and metrics. First, we show the case where simulations were performed for each of the five considered charge/discharge profiles (i.e. C/5, C/2, 1C, P and DST) as well as all possible combinations of these profiles, giving rise to the 31 profiles tested here. First, each of these 31 profiles are used in an optimization problem to find the optimal parameters for that particular profile by using a global minimizer as explained above. Thus, 31 optimal parameter sets in Eq.(14) are obtained, one per used profile, which is denoted as Cases. Each of these parameter vectors is then evaluated in terms of output voltage error, the RMSE in Eq.(15), for all the 31 considered profiles, which is denoted as Scenarios. A total of 961 validation test (31 parameter sets × 31 operating conditions) can be obtained. This yields the 3131 matrix shown in Figure 4 whose x-axis corresponds to the Scenarios (31 operating conditions) and the y-axis corresponds to the Cases (31 parameter sets). Each row of this matrix reflects a different parameter estimate for each considered profile, and each column is the model evaluation of that particular parameter vector considering the specific profile.
@@ 3.1 L0: Validation by Scenario 1
Analysis of both Figure 5 and Table 4 reveals that Case 1 required the highest total time among all Cases. The total time is calculated by Eq. (19). This extended duration is a direct consequence of its comprehensive inclusion of all five operating conditions, resulting in prolonged experimental procedures and parameter estimation processes. A particularly noteworthy observation is that Case 21{C/2,1C}’s parameter estimation error was second only to Case 1, while requiring merely 14% of Case 1’s total time. In other words, Case 21 achieved parameter estimation approximately seven times faster than Case 1.
@@ 3.2 L1: Overall Minimum RMSE Value
The L1 level is found by inspecting the minimum value(0.0073V) of the RMSE corresponding to the position (1,31) in the matrix using the (Case,Scenario) notation. In other words, use the first parameter set obtained from{C/5, C/2, 1C, P, DST} and verify it with the 31st operating condition {DST}. This tells us that the overall minimum error is obtained when the complete dataset is used to estimate the model parameters and only the DST data is used for model validation. Thus, the error obtained by considering the full dataset is larger than the one resulting from using DST alone (position (1,1) in the matrix), which implies that the error associated to the complementary part of the DST data in the full dataset (where full dataset is {C/5, C/2, 1C, P, DST} and the complement of the DST in the full dataset is {C/5, C/2, 1C, P}) involves a larger error than the DST itself. This statement is verified through the following derivation